Given a standard normal distribution with mean =61 and standard deviation = 35, find P(X<113)


1
Expert's answer
2022-02-04T13:36:05-0500

"\mu=61\\\sigma=35"

"p(x\lt 113)=p({x-\mu\over \sigma}\lt {113-\mu\over\sigma})=p(Z\lt{113-61\over35})=p(Z\lt1.49)=0.93119"